# Machine Unlearning Doesn’t Do What You Think: Lessons for Generative AI Policy, Research, and Practice

### v1

@@ 2.4 An expanded, loose definition for machine unlearning
In other words, the scope for unlearning no longer just concerns what we, following Cooper and Grimmelmann [26], will refer to as back-end considerations: “characteristics and capabilities of the model itself that directly result from its training.” Unlearning also concerns front-end considerations: “how the model behaves in generating outputs in response to … specific prompt[s]”(emphasis added) [26]. [^9] Data can enter a generative-AI system on the back-end as training data (i.e., for pre-training, fine-tuning, alignment) and the front-end via prompts, generation-time plug-ins, and retrieval-augmented generation (RAG). We refer to Lee et al. [77, Part I] for more information on data and the generative-AI supply chain. Both the back-end and front-end involve processes that have their own inputs and produce their own outputs (Figure 1). [^10] The utility of this framing becomes clear in Section 4, where we discuss concrete inputs and outputs of unlearning methods applied at these different stages. Using the words “input” and “output” would be unclear, as they are overloaded with different meanings at different stages. Also note that this is a different usage of “back-end” and “front-end” from Internet software, where “back-end” refers to server-side components like storage and “front-end” refers to client-side components like a user interface . On the back-end, the training dataset is an input and the trained model is an output; the back-end involves making choices for which training data to include, which training algorithm to run, etc. On the front-end, the prompt is an input and the generation is an output; the front-end includes the process of inference (i.e., producing generations), system-level filters that may prevent the processing of certain undesirable user prompts or the user-facing output of certain undesirable generations (Section 4.2), etc. On the back-end, the trained model is an output; on the front-end, the trained model is used to produce outputs. Throughout this paper, we will use this back-end/front-end terminology as a shorthand for distinguishing the different points in time where unlearning is of interest, and which artifacts a particular unlearning method is intended to affect—the model parameters or the model’s possible generations (Section 4).
Extending the example at the beginning of this section of an individual exercising their “right to be forgotten” under the GDPR: for a generative-AI model, the goal for unlearning would no longer simply be to remove the influence of the individual’s personal data from the model’s parameters on the back-end, but also to ensure that the resulting model could not produce outputs that reflect that individual’s personal data on the front-end. It is an appealing idea that machine unlearning could serve both of these ends. Indeed, it would be remarkably convenient if machine unlearning could be both an approach for mitigating the influence of problematic training data on a trained model’s parameters and for effectively moderating a generative-AI model’s possible outputs. If so, it would also perhaps be reasonable to assume, as many researchers and organizations have, that machine unlearning could, on its own, be used to solve issues related to problematic model outputs in a variety of policy-relevant domains: novel privacy challenges [73, 23, 12, 94, 148], copyright [43, 150, 77, 26, 137], safety [81, 86, 80], and more.
However, as we discuss below, these two back-end and front-end goals are very different in kind (Sections 45). Tying them together ultimately muddles what concrete unlearning methods could reasonably achieve on their own for desired policy ends (Section 6). To arrive at this conclusion, we first need some additional language that will enable us to refine our discussion of what types of information machine unlearning could address.
3 Targets for Machine Unlearning
@@
Our loose definition for unlearning is abstract (Section 2.4); in fact, it is so abstract that it allows for an enormous number of reasonable interpretations and possible techniques that satisfy it. In this section, our aim is to provide some language that can help us be more precise. Building on our loose definition, we now pin down useful ways to think about what a “piece of information” could mean—what types of information one might want to target with unlearning. In the sections that follow, we show that the targets we define are places where concrete unlearning methods could potentially apply in practice (Section 4). We will also note that some articulated goals for unlearning escape these target definitions altogether, highlighting instances where unlearning methods could not be applied rigorously or reliably for certain desired ends (Section 5).
@@
We define three overarching (and, as we will see, overlapping) targets: observed information (Definition 1), latent information (Definition 2), and higher-order concepts (Definition 3). These definitions build upon each other and become more abstract and indeterminate. After presenting the definitions, we discuss how this indeterminacy surfaces challenges for designing and implementing unlearning methods in practice.
@@
Observed information includes training-data examples: the contiguous pieces of data that are the base-level unit of input to model training (Section 2.1). For example, consider that the text “Susan’s phone number is 555-123-4567” is included as an example in an LLM’s training data. Since this text is used directly to train the LLM, it is observed information. Observed information also captures sets of training examples, such as all examples in the overall training dataset that mention Susan. It also includes data contained within examples, such as just the phone number “555-123-4567” in “Susan’s number is 555-123-4567.”
Effective trained models generalize: the learning process instills models with complex patterns that are derived from the observed information in the training data—patterns that models can apply to previously unseen information when they are put to use for inference or generation (Section 2.1). This learned information is latent in the training data.111111 For useful models, most learning is generalization; however, models also memorize (near) exactly a portion of their training data [e.g., 26, 77, 18, 108, 48, 121, 99]. Understanding the relationship between memorization and generalization is an active area of research.
@@
Unlike observed information (Definition 1), latent information is not literally observed in the training data. However, there are ML-based methods that claim to identify latent information and make it observable in the trained model’s parameters [^12] For example, some methods identify “concept neurons:” model parameters that relate to human-interpretable concepts [52, 40, 8]. or indirectly through a model’s outputs when the model is put to use. [^13] In some cases, it may be possible to identify non-exhaustively the data examples that contributed to latent information. However, currently, this is not true in general (Section 5). Eliciting information from the model’s parameters in useful ways is one of the goals of mechanistic interpretability research  [e.g., 98, 25, 97, 59]. Latent information can include simple deductions [68, 111]. For example, given the observed information “Carlos is going to Susan’s house for a birthday party this Thursday” and “Susan lives in Philadelphia,” a possible piece of latent information is that Carlos is going to be in Philadelphia on Thursday. [^14] Of course, just as with observed information, there is no guarantee that latent information is factually correct. (In this example, perhaps Carlos attends the party remotely over a video call.) This information is not literally contained in the training data; it is derived from relationships learned from observed information. Of course, latent information can also be significantly more complex than such simple deductions. The power of large-scale models trained on enormous datasets [76] comes from their flexibility to capture all sorts of latent information—across observed information, across latent information, or across some combination of the two. Indeed, information can interact to produce sophisticated, higher-order information that ML research often refers to as “knowledge” or “capabilities.”
@@
This spectrum reveals that Definition 3 is somewhat arbitrary, since it is not clear how to distinguish when a piece of latent information is sufficiently complex to be considered a higher-order concept. We do not attempt to draw these lines. Nevertheless, we still find it useful in the discussion that follows to have a target definition that lets us to refer to the unlearning of higher-order concepts, since this is a type of information that could reasonably be—and, in some cases, is claimed to be—a target for machine unlearning (Sections 456).

### v2

@@ Appendix B Additional notes on evolving motivations for unlearning
This back-end/front-end terminology can serve as useful shorthand for distinguishing the different points in time where unlearning is of interest, and which artifacts a particular unlearning method is intended to affect—the model parameters or the model’s possible generations (Section 3). Using the words “input” and “output” can sometimes be unclear, as they are overloaded with different meanings at different stages. (Also note that this is a different usage of “back-end” and “front-end” from Internet software, where “back-end” refers to server-side components like storage and “front-end” refers to client-side components like a user interface.)
Appendix C Targets for Machine Unlearning
@@
Our loose definition for unlearning is abstract (Section 2); in fact, it is so abstract that it allows for an enormous number of reasonable interpretations and possible techniques that satisfy it. In this appendix, our aim is to provide some language that can help us be more precise. Building on our loose definition, we pin down useful ways to think about what types of information one might want to target with unlearning. In the main paper, we informally reference these targets; they are places where concrete unlearning methods could potentially apply in practice (Section 3). When we discuss mistmatches (Section 4), we also note that some articulated goals for unlearning escape these target definitions altogether, highlighting instances where unlearning methods, from first principles, could not be applied rigorously or reliably for certain desired ends.
@@
We define three overarching (and, as we will see, overlapping) targets: observed information (Definition 1), latent information (Definition 2), and higher-order concepts (Definition 3). These definitions build upon each other and become more abstract and indeterminate. This indeterminacy surfaces challenges for designing and implementing unlearning methods (Sections 34).
@@
Observed information includes training-data examples. For instance, consider that the text “Susan’s phone number is 555-123-4567” is included as an example in an LLM’s training data. Since this text is used directly to train the LLM, it is observed information. Observed information also captures sets of training examples, such as all examples in the overall training dataset that mention Susan. It also includes data contained within examples, such as just the phone number “555-123-4567” in “Susan’s number is 555-123-4567.” Removal methods target observed information (Section 3.1).
Effective trained models generalize: the learning process instills models with complex patterns that are derived from the observed information in the training data—patterns that models can apply to previously unseen information when they are put to use for inference or generation. This learned information is latent in the training data. Removal methods target observed information directly (Section 3.1).
@@
Unlike observed information (Definition 1), latent information is not literally observed in the training data. (However, there are ML-based methods that claim to identify latent information and make it observable in the trained model’s parameters [e.g., 55, 42, 8] or indirectly through a model’s outputs when the model is put to use [e.g., 107, 25, 106, 63].)
Latent information can include simple deductions [73, 120]. For example, given the observed information “Carlos is going to Susan’s house for a birthday party this Thursday” and “Susan lives in Philadelphia,” a possible piece of latent information is that Carlos is going to be in Philadelphia on Thursday. (Of course, just as with observed information, there is no guarantee that latent information is factually correct. In this example, perhaps Carlos attends the party remotely over a video call.) This information is not literally contained in the training data; it is derived from relationships learned from observed information.
Latent information can also be significantly more complex than such simple deductions. The power of large-scale models trained on enormous datasets [85] comes from their flexibility to capture all sorts of latent information—across observed information, across latent information, or across some combination of the two. Indeed, information can interact to produce sophisticated, higher-order information that ML research often refers to as “knowledge” or “capabilities.”
@@
This spectrum reveals that Definition 3 is somewhat arbitrary, since it is not clear how to distinguish when a piece of latent information is sufficiently complex to be considered a higher-order concept. We do not attempt to draw these lines. Nevertheless, we still find it useful to have a target definition that lets us to refer to the unlearning of higher-order concepts, since this is a type of information that could reasonably be—and, in some cases, is claimed to be—a target for machine unlearning (Sections 345).